# Hallmarks of Optimization Trajectories in Neural Networks and LLMs: The Lengths, Bends, and Dead Ends

### v1

@@ Abstract
We propose a fresh take on understanding the mechanisms of neural networks by analyzing the rich structure of parameters contained within their optimization trajectories. Towards this end, we introduce some natural notions of the complexity of optimization trajectories, both qualitative and quantitative, which reveal the inherent nuance and interplay involved between various optimization choices, such as momentum, weight decay, and batch size. We use them to provide key hallmarks about the nature of optimization in deep neural networks: when it goes right, and when it finds itself in a dead end. Further, thanks to our trajectory perspective, we uncover an intertwined behaviour of momentum and weight decay that promotes directional exploration, as well as a directional regularization behaviour of some others. We perform experiments over large-scale vision and language settings, including large language models (LLMs) with up to 12 billion parameters, to demonstrate the value of our approach.
Machine Learning, ICML
[…]
@@ 1 Introduction
Given a network architecture and the training task, the loss landscape — which is the high-dimensional surface whose each point characterizes the fit of the parameters to the task objective — entails the possible trajectories that might be followed by an optimization algorithm, such as stochastic gradient descent (SGD). The particular sets of paths are determined, to no less extent, by the particular optimization choices and hyperparameters, such as the learning rate, momentum, batch size, weight decay, and more. We might even say that the regions and topographical features of the landscape that are never encountered or realized in typical optimization paths, might as well not be in the landscape. Essentially, the optimization trajectories are the probes through which the loss landscape is accessed.
@@
Consequently, a significant body of literature builds around the principle of an inherently regular manner of traversing in the landscape, i.e., the implicit bias (Gunasekar et al., 2018; Li et al., 2019, 2020; Moroshko et al., 2020), facilitated by optimization algorithms. This preferential landscape access is in comparison with the actual surface level, possibly treacherous, non-convexity, and thus suggesting that the network stays clear of sub-optimal local minima. Therefore, despite the complexity of the neural landscapes, implicit bias lends a formal and reasonable support to the empirical success of massively over-parameterized neural networks
If that is indeed the case, we might expect to see traits and hallmarks of regularity in the sequence of steps that make up the optimization trajectories of neural networks in the loss landscape. In other words, it raises the following questions,
@@
How are these trajectories structured? Are these paths long or short, with zigzags or soft bends, reaching the solution winding and coiling or straight and direct? What, if at all, do they reveal about the generalization strategy?
@@
This, in essence, is the key research theme of our study here. More precisely, we explore and develop key qualitative as well as quantitative indicators (hallmarks) about the complexity/regularity of the optimization trajectory. In particular, towards this end, we analyze and compare multiple intermediate checkpoints amongst themselves, across different scenarios and large-scale case studies. The qualitative hallmark, which we call the ‘Trajectory Map’, conveys the directional (dis)similarity of the parameters and visually depicts the nature of optimization within and across various stages of training, i.e. at a pan-trajectory level. The quantitative hallmarks are functions of these trajectory maps, that measure various notions of lengths and angles, over and about the sequence of steps in the trajectory.
@@
We form it as a focus of our investigation to study the properties of trajectories and their ensuing implications about the nature of found solutions since it, (a) brings in a level of architecture agnosticity and helps unlock shared insights onto features of optimization, (b) contains an intrinsic data-dependence that, for all intents and purposes, implies no explicit inference over additional data samples is needed (which anyway might not be possible due to resource or privacy constraints), (c) allows analyzing and prognosing the developing generalization strategy on-the-fly, over the course of training (instead of waiting all the way until convergence to assess the solution quality), and (d) provides potential hints at the bottlenecks and redundancies plaguing the optimization procedure.
Hence, given these factors, we naturally do not expect the above trajectory viewpoint to cover and uncover all aspects of the desired phenomenology at stake, but we hope to at least offer an outlook into the concerned phenomenon based on the corresponding optimization trajectories (and their directional complexity), such as when considering the implicit effects of momentum and weight decay, and thus provide novel insights into the inner-workings of neural networks and their optimization.
@@ 2 Methodology
Let us assume the optimization trajectory consists of a set $\mathcal{T}$ of points $\{{\boldsymbol{\theta}}_{t}\}_{t=0}^{T}$, each denoting the (flattened) parameters of the network encountered at some step and which live in the parameter space $\Theta={\mathbb{R}^{p}}$, i.e., $\mathcal{T}\subseteq\Theta$. This set of points need not contain the entire set of points visited in the course of optimization but instead can represent a subset of points, possibly sampled at an interval of $k$ points. [^1] In practice, such sub-sampling will indeed be required to avoid extremely high storage costs to checkpoint the parameters for each point, and the sub-sampling interval will be typically in epochs. It will be convenient to organize this set of points which define the trajectory in the form of a matrix, ${\boldsymbol{\Theta}}\in{\mathbb{R}^{(T+1)\times p}}$, whose first dimension $T+1$ makes explicit the inclusion of the initialization ${\boldsymbol{\theta}}_{0}$.
@@
Elsewhere it will be useful to analyze the trajectory relative to some point ${\boldsymbol{\theta}}_{\tau}$ as the origin, i.e., we will instead consider the set of points $\mathcal{T}_{\tau}=\{{\boldsymbol{\theta}}_{t}-{\boldsymbol{\theta}}_{\tau}\}_{t=0}^{T}$, and correspondingly organize it in the matrix ${\boldsymbol{\Theta}}_{\tau}\in{\mathbb{R}^{(T+1)\times p}}$. When $\tau$ is itself one of the points of the trajectory, then we will omit the row of zeros and shape the matrix of size ${\mathbb{R}^{T\times p}}$. A natural point from where to view the trajectory would be the initialization ${\boldsymbol{\theta}}_{0}$, and this relative trajectory will then be denoted as ${{\boldsymbol{\Theta}}}_{0}$ (where the subscript 00 is not to be confused for the usual origin $O$, namely, ${\boldsymbol{\theta}}_{O}=\mathbf{0}$).
Trajectory Map.
@@
Analyzing the matrix ${\boldsymbol{\Theta}}$ or ${\boldsymbol{\Theta}}_{\tau}$, on its own, might get cumbersome as the size of modern networks ranges in millions and billions of parameters. Hence, we will resort to looking at functions of the kernel matrix ${\bf K}={\boldsymbol{\Theta}}{\boldsymbol{\Theta}}^{\top}$ which would be a square matrix of shape $n=T+1$, or for $\tau\neq O$, the relative kernel matrix ${\bf K}_{\tau}={\boldsymbol{\Theta}}_{\tau}{\boldsymbol{\Theta}}_{\tau}^{\top}$ of shape $n=T+1$ or $n=T$ depending if the point ${\boldsymbol{\theta}}_{\tau}$ is a part of the trajectory or not. Further, it will also be helpful to isolate and analyze the directional aspect of the trajectory, for which we will normalize the set of points by their norm, and in effect, consider the set $\widehat{\mathcal{T}}_{\tau}=\{\frac{{\boldsymbol{\theta}}_{t}-{\boldsymbol{\theta}}_{\tau}}{\|{\boldsymbol{\theta}}_{t}-{\boldsymbol{\theta}}_{\tau}\|_{2}}\}_{t=0}^{T}$ with the respective matrix $\widehat{{\boldsymbol{\Theta}}}_{\tau}$. As a result, the ensuing kernel matrix $\widehat{{\boldsymbol{\Theta}}}_{\tau}\widehat{{\boldsymbol{\Theta}}}_{\tau}^{\top}$, which we will refer to as ${\bf C}_{\tau}$ (or ${\bf C}:={\bf C}_{O}$ for the usual origin $\tau=O$), will contain the relative cosine similarities between every pair of points in the trajectory. So, $({\bf C}_{\tau})_{ij}$ is,
@@
Quantitative Hallmarks.
Besides visualizing the TM as the qualitative hallmark, we will consider the following set of indicators for quantitatively hallmarking the optimization trajectories.
@@
To get a better sense of MDS, we can consider its two possible extremes: (a) all the parameter unit-vectors cancel out, yielding a value of $\omega=0$. For instance, this would happen in the scenario when the points in the trajectory are exactly following a circular orbit around the origin; or, (b) when each of the parameters point in the same direction, implying that the trajectory is simply a linear path, with $\omega=1$. Knowing the nature of these two extremes, we can expect neither of them to be desirable in an ideal trajectory which leads up to a generalizing solution. Thus, hitting its sweet spot would be the target, or where that is unknown, at least avoiding these extremes.
@@
(iii) Norm-based Measures: Besides, we will also include other previously reported measures of lengths that are relevant in this respect, such as: parameter norms $\|{\boldsymbol{\theta}}_{t}\|_{2}$, distance from initialization $\|{\boldsymbol{\theta}}_{t}-{\boldsymbol{\theta}}_{0}\|_{2}$, norm between consecutive points $\|{\boldsymbol{\theta}}_{t+1}-{\boldsymbol{\theta}}_{t}\|_{2}$.
Remark. For extremely large neural networks, with several billions of parameters, building the underlying kernel matrices can start being a tad bit resource-expensive. In principle, there is a rich body of work in kernel methods that has focused on developing efficient approximations (Davis et al., 2014; Chen et al., 2021). But, as we foray into this novel trajectory perspective, we do not resort to such approximations and compute measures exactly for the sake of accuracy.
@@ 3 A Tale of Hyperparameters
There is an emerging folk intuition in the research community that, over the past decade, networks have co-evolved hand-in-hand, amidst other things, with a certain particular set of optimizers and hyperparameter choices, deviating from which tends to produce significantly poor results. Let us see one such case in action, for the case of ResNet50 trained on ImageNet with SGD, achieving the familiar top-1 accuracy of $\sim 76\%$. Further, as is usual convention, this network was trained via a learning rate $\eta=0.1$, momentum $\mu=0.9$, batch size $B=256$, weight decay $\lambda=0.0001$ for $90$ epochs, with a multiplicative decay by a factor of $0.1$ at epochs $30$ and $60$.
@@
Before we start this excursion with the hyperparameters, let us simply pay a glance at the nature of the optimization trajectory, as visualized via the proposed qualitative hallmark, namely the trajectory map ${\bf C}$. In particular, we save checkpoints from the initialization until the final epoch, at intervals of every epoch, and thereby giving us a total of $91$ checkpoints. We plot this $91\times 91$ matrix in Figure 2.
For starters, we can easily make out three [^2] If we look closely, there seems to be another phase transition neighbouring the initialization and the subsequent couple of epochs, giving rise to a thin horizontal and vertical sliver of relatively lighter colour in this figure. distinct phases of optimization which is marked by an increased darkening of the pixels and their locations are precisely where learning rate decay was applied. The onset of these phases also seems to bring about an increased cosine similarity of the parameters contained within these phases, which seems to imply that following the learning rate decay the optimization is honing into a progressively confined subspace of the landscape.
@@
Next, the mean directionality score (MDS) for this particular figure comes out to be $\omega=0.764$. To contextualize this value, we should remind ourselves that we are working in a space of $\sim 25.6$ million dimensions. This should serve to emphasize that network optimization trajectories are highly structured, and not merely random points in high-dimensions whose cosine similarity goes to zero. As a further note, the cosine similarity of two different instantiations of the ResNet50 parameters, from the same (and the usual) random initialization scheme, gives a value [^3] A reason for this high value of cosine similarity is the presence of BatchNorm or LayerNorm learnable scale parameters, which are typically initialized to all ones. of $0.374$. However, besides these comparisons, it is still somewhat unclear what the value of MDS tells us about the trajectory. In particular, we need to ask ourselves (i) How does it rank in comparison to possibly other trajectories? (ii) Is it a high enough value, or should we aim for something much larger? (iii) But since the maximum value of $\omega$ can be $1$ and since we have initialization in the mix here, obtaining such a value would not be so desirable as it would effectively mean the absence of any feature learning (Chizat et al., 2020), which seems to be a critical component behind the success of deep learning.
@@
To better address these and other burning questions, let us now tinker slightly with certain hyperparameters. In particular, we will only touch just two of them, namely, momentum and weight decay, and turn them off one at a time. Importantly, everything else remains fixed. The ensuing trajectory maps can be found in Figure 3, where either momentum is turned off ($\mu=0$) or weight decay ($\lambda=0$) or both ($\mu=0,\lambda=0$). The first thing that stands out is that the trajectory maps become darker in colour, suggesting an increased cosine similarity, and especially the subfigure without both momentum and weight decay looks rather conspicuous, with essentially just one dark grid of extremely large cosine similarities. In particular, if we calculate the mean directional similarities, they turn out to be $\omega=0.901,0.931,0.979$, respectively. It would also be worth noting that their top-1 accuracies are, $72.63\%,70.86\%,68.73\%$, with a net drop of about $10\%$ by just turning off these two hyperparameters.
These results would indicate that the optimization hyperparameters seem to be leading to qualitatively different solutions, with the presence of weight decay or momentum encouraging more directional exploration, whereas in their absence, the optimization latches on to a nearby solution (at least, in a directional sense) and then short-circuits itself into a ‘dead end’. However, these results also seem to run against the simplistic pictures one might conceive of regarding momentum and weight decay. For instance, intuitively, the use of momentum should add strength to the previous gradient directions and lead to increased directional similarity, but we find the opposite to hold. Similarly, in the absence of weight decay, the network is not constrained to remain in a ball around the origin and, in principle, there should be more license to explore in the landscape. Or, what are these intuitive notions not taking into account?

### v2

# Hallmarks of Optimization Trajectories in Neural Networks: Directional Exploration and Redundancy

@@ Abstract
We propose a fresh take on understanding the mechanisms of neural networks by analyzing the rich directional structure of optimization trajectories, represented by their pointwise parameters. Towards this end, we introduce some natural notions of the complexity of optimization trajectories, both qualitative and quantitative, which hallmark the directional nature of optimization in neural networks: when is there redundancy, and when exploration. We use them to reveal the inherent nuance and interplay involved between various optimization choices, such as momentum and weight decay. Further, the trajectory perspective helps us see the effect of scale on regularizing the directional nature of trajectories, and as a by-product, we also observe an intriguing heterogeneity of Q,K,V dynamics in the middle attention layers in LLMs and which is homogenized by scale. Importantly, we put the significant directional redundancy observed to the test by demonstrating that training only scalar batchnorm parameters some while into training matches the performance of training the entire network, which thus exhibits the potential of hybrid optimization schemes that are geared towards efficiency.
@@ 1 Introduction
Given a network architecture and the training task, the loss landscape — which is the high-dimensional surface whose each point characterizes the fit of the parameters to the task objective — entails the possible trajectories (or paths) that might be followed by an optimization algorithm, such as stochastic gradient descent (SGD). The particular sets of optimization trajectories are determined, to no lesser extent, by the particular optimization choices and hyperparameters, such as the learning rate, momentum, batch size, weight decay, and more. We might even say that the regions and topographical features of the landscape that are never encountered or realized in typical optimization trajectories, might as well not be in the landscape at all. Essentially, the optimization trajectories are the probes through which the loss landscape is accessed.
@@
Consequently, a significant body of literature builds around the principle of an inherently regular manner of traversing in the landscape, i.e., the implicit bias [Gunasekar et al., 2018; Li et al., 2019, 2020; Moroshko et al., 2020], facilitated by optimization algorithms. This preferential landscape access is in the face of the actual surface level, possibly treacherous, non-convexity, and thus suggesting that the network stays clear of sub-optimal local minima. Therefore, despite the complexity of the neural landscapes, implicit bias lends a formal and reasonable support to the empirical success of massively over-parameterized neural networks. If that is indeed the case, we might expect to see traits and hallmarks of regularity in the sequence of steps that make up the optimization trajectories of neural networks in the loss landscape. In other words, it provokes the following questions,
@@
How are these trajectories structured? Do these paths have a lot of zigzags and bends, reaching the solution winding and coiling, or are they straight and direct? And does this depend upon the phase of optimization (early vs late)?
@@
This, in essence, is the key research theme of our present study. More precisely, we explore and develop key qualitative as well as quantitative indicators (hallmarks) about the complexity/regularity of the optimization trajectory. Towards this end, we analyze and compare multiple intermediate checkpoints amongst themselves, across different scenarios and large-scale case studies. A qualitative hallmark, which we call the ‘Trajectory Map’, conveys the directional (dis)similarity of the parameters and visually depicts the nature of optimization within and across various stages of training, i.e., at a pan-trajectory level. Our quantitative hallmarks are functions of these trajectory maps, measuring various notions of lengths and angles, over and about the sequence of steps in the trajectory.
@@
The focus of our investigation is to study the properties of trajectories and their ensuing implications about the nature of found solutions since it (a) brings in a level of architecture agnosticity and helps unlock shared insights onto features of optimization, (b) contains an intrinsic data-dependence that, for all intents and purposes, necessitates no explicit inference over additional data samples is needed (which anyway might not be possible due to resource or privacy constraints), (c) allows analyzing and prognosing the developing solution strategy on-the-fly, over the course of training (instead of waiting all the way until convergence to assess the solution quality), and (d) provides potential hints at the bottlenecks and redundancies plaguing the optimization procedure.
Our contributions are: 1) We propose the novel perspective of trajectory maps and showcase its use for hallmarking the directional nature of optimization. 2) We utilize it to theoretically show the intertwined effects of momentum and weight decay on obtuse angles between updates. 3) We show how scale provides a regularizing effect on the directional complexity of trajectories, and also find an interesting observation about Q,K,V heterogeneity in the middle attention layers. 4) We put the demonstrated directional redundancy to the test, by showing that only training the scalar parameters of (batch) normalization layers a short while into training suffices for matching the performance, thus laying the seeds for efficient optimization hybrids.
@@ 2 Methodology
Matrix representation of Trajectory. Let us assume the optimization trajectory consists of a set $\mathcal{T}$ of points $\{{\boldsymbol{\theta}}_{t}\}_{t=0}^{T}$, each denoting the (flattened) parameters of the network encountered at some step and which live in the parameter space ${\boldsymbol{\theta}}={\mathbb{R}^{p}}$, i.e., $\mathcal{T}\subseteq{\boldsymbol{\theta}}$. This set of points need not contain the entire set of points visited in the course of optimization but instead can represent a subset of points, possibly sampled at an interval of $k$ points. It will be convenient to organize this set of points, which define the trajectory, in the form of a matrix, ${\boldsymbol{\Theta}}\in{\mathbb{R}^{(T+1)\times p}}$, whose first dimension $T+1$ makes explicit the inclusion of the initialization ${\boldsymbol{\theta}}_{0}$.
@@
Elsewhere it might be useful to analyze the trajectory relative to some point ${\boldsymbol{\theta}}_{\tau}$ as the origin, so there we will instead consider the set of points $\mathcal{T}_{\tau}=\{{\boldsymbol{\theta}}_{t}-{\boldsymbol{\theta}}_{\tau}\}_{t=0}^{T}$, and correspondingly organize it in the matrix ${\boldsymbol{\Theta}}_{\tau}\in{\mathbb{R}^{(T+1)\times p}}$. When $\tau$ is itself one of the points of the trajectory, then we will omit the row of zeros and shape the matrix as ${\mathbb{R}^{T\times p}}$. A natural point from where to contextualize the trajectory would be the initialization ${\boldsymbol{\theta}}_{0}$, and this relative trajectory will then be denoted as ${{\boldsymbol{\Theta}}}_{0}$ (where the subscript 00 is not to be confused for the usual origin $O$, namely, ${\boldsymbol{\theta}}_{O}=\mathbf{0}$).
@@
Trajectory Map. Analyzing the matrix ${\boldsymbol{\Theta}}$ or ${\boldsymbol{\Theta}}_{\tau}$, on its own, might get cumbersome as the size of modern networks ranges in millions and billions of parameters. Hence, we will resort to looking at functions of the kernel matrix ${\bf K}={\boldsymbol{\Theta}}{\boldsymbol{\Theta}}^{\top}$ which would be a square matrix of shape $n=T+1$, or for $\tau\neq O$, the relative kernel matrix ${\bf K}_{\tau}={\boldsymbol{\Theta}}_{\tau}{\boldsymbol{\Theta}}_{\tau}^{\top}$ of shape $n=T+1$ or $n=T$ depending if the point ${\boldsymbol{\theta}}_{\tau}$ is a part of the trajectory or not. Further, it will also be helpful to isolate and analyze the directional aspect of the trajectory, for which we will normalize the set of points by their norm, and in effect, consider the set $\widehat{\mathcal{T}}_{\tau}=\{\frac{{\boldsymbol{\theta}}_{t}-{\boldsymbol{\theta}}_{\tau}}{\|{\boldsymbol{\theta}}_{t}-{\boldsymbol{\theta}}_{\tau}\|_{2}}\}_{t=0}^{T}$ with the respective matrix $\widehat{{\boldsymbol{\Theta}}}_{\tau}$. As a result, the ensuing kernel matrix $\widehat{{\boldsymbol{\Theta}}}_{\tau}\widehat{{\boldsymbol{\Theta}}}_{\tau}^{\top}$, which we will refer to as ${\bf C}_{\tau}$ (or ${\bf C}:={\bf C}_{O}$ for the usual origin $\tau=O$), will contain the relative cosine similarities between every pair of points in the trajectory. So, $({\bf C}_{\tau})_{ij}$ is,
@@
Quantitative Hallmarks. Besides visualizing the TM as the qualitative hallmark, we will consider the following set of indicators for quantitatively hallmarking the optimization trajectories.
@@
To get a better sense of MDS, we can consider its two possible extremes: (a) all the parameter unit-vectors cancel out, yielding a value of $\omega=0$. For instance, this would happen in the scenario when the points in the trajectory are exactly following a circular orbit around the origin; or, (b) when each of the parameters point in the same direction, implying that the trajectory is simply a linear path, with $\omega=1$. Knowing the nature of these two extremes, we can expect neither to be desirable in an ideal trajectory which leads up to a generalizing solution. Thus, hitting its sweet spot would be the target, or where that is unknown, at least avoiding these extremes.
@@
(iii) Norm-based Measures: Besides, we also include common measures such as: parameter norms $\|{\boldsymbol{\theta}}_{t}\|_{2}$, distance from initialization $\|{\boldsymbol{\theta}}_{t}-{\boldsymbol{\theta}}_{0}\|_{2}$, norm between consecutive points $\|{\boldsymbol{\theta}}_{t+1}-{\boldsymbol{\theta}}_{t}\|_{2}$.
Remark. For extremely large neural networks, building the underlying kernel matrices can start being a tad bit resource-expensive. In principle, there is a rich body of work in kernel methods that has focused on developing efficient approximations [Davis et al., 2014; Chen et al., 2021]. But, as we foray into this novel trajectory perspective, we do not resort to such approximations for the sake of accuracy.
@@ 3 A Tale of Hyperparameters
There is an emerging folk intuition that, over the past decade, networks have co-evolved hand-in-hand, amidst other things, with a certain particular set of optimizers and hyperparameter choices, deviating from which tends to produce significantly poor results. Let us see one such case in action, for a ResNet50 trained on ImageNet with SGD, achieving the familiar top-1 accuracy of $\sim 76\%$. Further, as is usual convention, this network was trained via a learning rate $\eta=0.1$, momentum $\mu=0.9$, batch size $B=256$, weight decay $\lambda=0.0001$ for $90$ epochs, with a multiplicative decay by a factor of $0.1$ at epochs $30$ and $60$.
@@
Before we start the hyperparameter excursion, let us look at the nature of the optimization trajectory, as visualized via the proposed qualitative hallmark, namely the trajectory map ${\bf C}$. In particular, we save epochly checkpoints from the initialization until the final epoch, and thereby giving us a total of $91$ checkpoints. We plot this $91\times 91$ matrix in Figure 2.
For starters, we can easily make out three [^1] If we look closely, there seems to be another phase transition neighbouring the initialization and the subsequent couple of epochs, giving rise to a thin horizontal and vertical sliver of relatively lighter colour in this figure. distinct phases of optimization which are marked by an increased darkening of the pixels and their locations are precisely where learning rate decay was applied. The onset of these phases also seems to bring about an increased cosine similarity of the parameters contained within these phases, which seems to imply that following the learning rate decay the optimization is honing into a progressively confined subspace of the landscape.
@@
Next, the mean directionality score (MDS) for this particular figure comes out to be $\omega=0.764$. To contextualize this value, we should remind ourselves that we are working in a space of $\sim 25.6$ million dimensions. This should serve to emphasize that network optimization trajectories are highly structured, and not merely random points [^2] In Section C, we analyze the relative trajectory map and MDS for a random walk/Brownian motion, and in comparison we find that (expectedly) the trajectory maps of neural networks are more directionally redundant. in high-dimensions whose cosine similarity goes to zero. As a further note, the cosine similarity of two different instantiations of the ResNet50 parameters, from the same (and the usual) random initialization scheme, gives a value [^3] A reason for this high value of cosine similarity is the presence of BatchNorm or LayerNorm learnable scale parameters, which are typically initialized to all ones. of $0.374$. However, besides these comparisons, it is still somewhat unclear what the value of MDS tells us about the trajectory. In particular, we need to ask ourselves (i) How does it rank in comparison to possibly other trajectories? (ii) Is it a high enough value, or should we aim for something much larger? (iii) But since the maximum value of $\omega$ can be $1$ and since we have initialization in the mix here, obtaining such a value would not be so desirable as it would effectively mean the absence of any feature learning [Chizat et al., 2020], which seems to be a critical component [^4] Also, a certain amount of directional exploration is crucial as evident from our analysis of the grokking [Power et al., 2022] phenomenon discussed in Section H. behind deep learning’s success.
@@
To better address these questions, let us now tinker slightly with certain hyperparameters. In particular, we will only touch just two of them, namely, momentum and weight decay, and turn them off one at a time. Importantly, everything else remains fixed. The ensuing trajectory maps can be found in Figure 3, where either momentum is turned off ($\mu=0$) or weight decay ($\lambda=0$) or both ($\mu=0,\lambda=0$). The first thing that stands out is that the trajectory maps become darker in colour, suggesting an increased cosine similarity, and especially the subfigure without both momentum and weight decay looks rather conspicuous, with essentially just one dark grid of extremely large cosine similarities. In particular, MDS here is $\omega=0.901,0.931,0.979$, respectively. This toggling of the hyperparameters not only has a significant effect on the trajectory maps but also their top-1 accuracies, which are $72.63\%,70.86\%,68.73\%$ respectively.
These results would indicate that the optimization hyperparameters seem to be leading to qualitatively different solutions, with the presence of weight decay or momentum encouraging more directional exploration, whereas in their absence, the optimization latches on to a nearby solution (at least, in a directional sense). However, these results also seem to run against the simplistic pictures one might conceive of momentum and weight decay. For instance, intuitively, the use of momentum should add strength to the previous gradient directions and lead to increased directional similarity, but we find the opposite to hold. Similarly, in the absence of weight decay, the network is not constrained to remain in a ball around the origin and, in principle, there should be more license to explore in the landscape. Or, what are these intuitive notions not taking into account?
4 A Ride with Momentum & Decay
In order to obtain a more refined understanding of the above observation, we would like to inspect in detail the angular and norm-based measures of the trajectory. We will largely [^5] In the Appendix E, we also present a wider set of results, namely, the directional effects of hyper-parameters such as learning rate and batch size, recent regularizers like Sharpness-Aware Minimization [Foret et al., 2021] (SAM), as well as more datasets and architectures such as Vision Tranformers on ImageNet and VGG16 on CIFAR10, and for different amounts of label noise in the dataset. focus on momentum and weight decay, given the especially intriguing directional nature of the trajectories caused by them. Now, let us start by taking a look at how aligned are the net or the aggregate updates [^6] We qualify this by ‘net’ or ‘aggregate’ as we are working on somewhat coarser granularity (i.e., $1$ epoch) than every update or step. But our current granularity is still rich enough to allow for the presented trends to persist even if go 2×2$\times$2 × to 5×5$\times$5 × more coarser. in the presence of these hyperparameters and otherwise.